OpenStudy (anonymous):

Solve: 4(x − 1) − (x + 6) = 2

OpenStudy (anonymous):

Use the distributive property.

OpenStudy (anonymous):

Still Confused

OpenStudy (anonymous):

Have you studied the distributive property before? I know I posted it on an earlier question.

OpenStudy (anonymous):

4(x − 1) − (x + 6) = 2 4x-4 -x-6 = 2 4x-x -4-6 = 2 3x- 10 = 2 3x = 2+10 x = 12/3 x = 4 Medal?

OpenStudy (anonymous):

@darkbbn I noticed that you're asking a lot of questions on the same properties. Are you wanting to learn the material, or have others do the work for you ? @Nancy_Lam please don't give out the answers. This site is here to help people learn, not just give them the answer.